At a particular store they sell coffee for $4 a pound, cardamon for $35 a pound, and a coffee and cardamon mixture for $5 a pound. Assuming no change in profit on selling the mix versus its components how much caradamon should be mixed with 10 pounds of coffee?
Answer by checkley75(3666)

4*10+35*X=5(10+X)
40+35X=50+5X
35X-5X=50-40
30X=10
X=10/30
X=1/3 POUND OF $35COMPONENTS SHOULDBE USED.
PROOF
4*10+35*(1/3)=5(10+1/3)
40+35/3=50+5/3
40+11 2/3=50+1 2/3
51 2/3=51 2/3